A 22-year-old man was in a motorcycle accident with resultant neck injuries that led to partial paralysis of the upper and lower limbs. Almost immediately his chest felt heavy and he became dyspneic. His pulmonary function values were as Vital capacity (supine) 650 mL Minute ventilation (supine) 6 L/min Respiratory rate (supine) 30 b/min PaO2 61 mm Hg PaCO2 47 mm Hg 1. What are the values for the tidal volume (TV) and the alveolar ventilation (AV) for this individual? (Assume a normal value for dead space.) Compare with normal. 2. What is contributing to the decreased alveolar ventilation? 3. Which nerves and muscles are involved in this problem? 4. How is the residual volume affected? 5. Define dyspnea 6. What is causing the dyspnea in this individual? 7. Describe the reflexes involved in causing tachypnea in this individual
Answer #1

From the above data,
Minute Ventilation (MV)= 6L/min
Respiratory Rate(RR) = 30 b/min
1.) Minute Ventilation (MV) = Tidal Volume(TV) X Respiratory Rate(RR)
TV= (MV)/(RR)
TV = 6/30 =0.2 litres=200 ml
Alveolar ventilation (AV) = [tidal volume − dead space] × respiratory rate =[200 ml- 150ml] x 30 =50 X 30 =1500ml/min.
The normal value for dead space is 150 ml.
Comparison with normal values:
The normal value for TV is 500 ml whereas the individual have a TV of 200 ml which is quite deficient.
The normal value for Alveolar ventilation (AV)is 4200 ml/min whereas the individual have a AV of 1500 ml/min which is quite low as compared to normal.

2.) The decreased alveolar ventilation is probably due to increase in PaCO2.The normal value of PaCO2 is 38-42 mmHg but in this case it is elevated,47 mmHg.

3.) The muscles of the lower rib cage and thoracic cavity is involved.The nerves of the spinal cord are affected.

4.) The residual volume is decreased.

5.) Dyspnea is difficult breathing or when one has to put labor to breathe or shortness of breath.

6.)The tidal volume is lowered by the reduced diaphragmatic and intercoastal innervation and this results in dyspnea in this individual.

7.) The reflexes involved in causing tachypnea in this individual are the central and peripheral chemoreceptors. The decreased PaO2 of 61 mmHg in the individual is sensed by the peripheral chemoreceptors and medulla sends out signal and as a result the rate of breathing is increased.The increase in PaCO2 is sensed by the central chemoreceptors.

.